OpenStudy (anonymous):

a/b+c/d=b/a+d/c,then prove that a3/b3+c3/d3=b3/a3+d3+c3.

OpenStudy (nikvist):

\[\frac{a}{b}+\frac{c}{d}=\frac{b}{a}+\frac{d}{c}\quad\Rightarrow\quad \frac{a^3}{b^3}+\frac{c^3}{d^3}=\frac{b^3}{a^3}+\frac{d^3}{c^3}\] Is this correct?